# Real-Time Video Generation with Pyramid Attention Broadcast

### v1

@@ 3.4 Scaling ability
To evaluate our method’s scalability, we conduct a series of experiments. In each experiment, we apply PAB246 (the higest quality, but less speedup) to Open-Sora as our baseline configuration, change only the video sizes, parallel method and GPU numbers.
Table 4: Communication and latency comparison of different sequence parallelism methods on 8 NVIDIA H100 GPUs with and without our method. original refers to our method on single GPU. comm. represents the total communication volume required to generate a single 2s 480p video.
@@
We compare the scaling efficiency with and without our method using 8 GPUs in Table 3.4 for four sequence parallelism methods including Megatron-SP (Korthikanti et al., 2023), DS-Ulysses (Jacobs et al., 2023) and DSP (Zhao et al., 2024). The results demonstrate that: i) PAB significantly reduces communication volume for all sequence parallelism methods. Furthermore, our method achieves the lowest communication cost compared to other techniques, and achieving near-linear scaling on 8 GPUs. With a larger temporal broadcast range, it can yield even greater performance improvements. ii) Implementing sequence parallelism alone is insufficient for optimal performance because of the significant communication overhead across multiple devices.
@@
We evaluate our method’s speed in terms of FPS on 8 and 16 devices. Since in inference, the batch size of diffusion is often 2 because of CFG. Therefore, we split the batch first and apply sequence parallelism to each batch; in this way, PAB can extend to 16 devices with almost linear acceleration. As shown in Figure 9, we can achieve real-time with very high FPS video generation for 480p videos on 8 devices and even for 720p on 16 devices. Note that with acceleration techniques like Torch Compiler, we are able to achieve even better speed.
@@
To further investigate how our method achieves such significant speedup, we provide a breakdown of the time consumption for various components, as shown in Figure 11. The analysis reveals that the attention calculation itself does not consume a large portion of time because the sequence length for attention will be much shorter if we do attention seperately for each dimension. However, attention-related operations, such as normalization and projection, are considerably more time-consuming than the attention mechanism itself, which mainly contribute to our speedup.
@@ 3.5 Visualization
As shown in Figure 12, we visualize the video results generated by our method compared to the original model. The generation specifications are the same with those in Table 1, and we employ the highest quality strategy outlined in the table. The visualized results demonstrate that our method maintains the original quality and details.
@@ 4.1 Video Generation
Video generation has seen remarkable progress recently, which aims to synthesize visually high-quality and motion-consistent videos. Previous research can be broadly divided into three key stages: GAN-based, auto-regressive, and diffusion models. Initially, GANs were extended from image to video generation, aiming to maintain spatiotemporal coherence (Vondrick et al., 2016; Saito et al., 2017; Clark et al., 2019; Kahembwe & Ramamoorthy, 2020; Wang et al., 2020). Despite their success, GAN-based approaches often struggled with mode collapse, which hindered their effectiveness in producing diverse and consistent videos. To address these limitations, auto-regressive models were introduced, leveraging Transformer architectures to predict video frames sequentially, conditioned on previously generated frames.  (Rakhimov et al., 2020; Weissenborn et al., 2020; Yan et al., 2021; Ge et al., 2022) Although these models generally yield higher video quality and more stable convergence, they demand substantial computational resources.
The latest advancements in video generation have been driven by diffusion models, which iteratively refine noisy inputs to generate high-fidelity video frames (Ho et al., 2022b; An et al., 2023; Esser et al., 2023; Ge et al., 2023). While many works focus on conv-based diffusions (Wang et al., 2020; Ho et al., 2020; Harvey et al., 2022; Mei & Patel, 2022; Singer et al., 2022; Ho et al., 2022a; Luo et al., 2023; Wang et al., 2023b; Wu et al., 2023; Zhang et al., 2023), researchers begin to explore Transformer-based diffusion models for video generation (Zheng et al., 2024; Lab & etc., 2024; Ma et al., 2024a) because of scalability (Peebles & Xie, 2023) and efficiency. In this work, we focus on accelerating the DiT-based video generation models.
@@ 4.2 Diffusion Model Acceleration
Advancements in video diffusion models have demonstrated their potential for high-quality video generation, yet their practical application is often limited by slow inference speeds. Previous research about speeding up diffusion model inference can be broadly classified into three categories. First, reducing the sampling time steps has been explored through methods such as DDIM (Song et al., 2020), which enables fewer sampling steps without compromising generation quality. Other works also explore efficient solver of ODE or SDE (Song et al., 2021; Jolicoeur-Martineau et al., 2021; Lu et al., 2022; Karras et al., 2022; Lu et al., 2023), which employs a pseudo numerical method to achieve faster sampling. Second, researchers aimed at reducing the workload and inference time at each sampling step, including distillation (Salimans & Ho, 2022; Li et al., 2023d), quantization (Li et al., 2023c; He et al., 2023; So et al., 2023a; Shang et al., 2023), distributed inference (Li et al., 2024). Third, jointly optimized methods simultaneously optimize network and sampling methods (Li et al., 2023a; Liu et al., 2023). Moreover, researchers modify the model structure indirectly by using the cache mechanism to reduce computation. Cache (Smith, 1982) in computer systems is a method to temporarily store frequently accessed data from the main memory to improve processing speed and efficiency. Based on the findings that high-level features usually change minimally between consecutive steps, researchers reuse the high-level features in U-Net structure while updating the low-level ones  (Ma et al., 2024c; Li et al., 2023b; Wimbauer et al., 2024; So et al., 2023b). Besides,  (Zhang et al., 2024) cache the redundant cross-attention in the fidelity-improving stage.
However, previous methods mainly focus on the U-Net structure and image domain. The most similar work are  (Ma et al., 2024b),  (Chen et al., 2024b),  (Zhang et al., 2024) and  (Li et al., 2024).  (Ma et al., 2024b) skip the computation of a large proportion of feedforward layers in DiT models through post-training.  (Zhang et al., 2024) cache the self-attention in the initial stage and reuses cross-attention in the fidelity-improving phase.  (Chen et al., 2024b) caches feature offsets of DiT blocks.  (Li et al., 2024) reduce the latency of single-sample generation by running convolution-based diffusion models across multiple devices in parallel while sacrificing quality and efficiency for parallel. Different from previous works, we aim at real-time DiT-based video generation models using training-free acceleration methods. We utilize pyramid attention and broadcast sequence parallel to accelerate video generation without loss of quality.
@@ 5 Discussion and Conclusion
In this work, we introduce Pyramid Attention Broadcast (PAB), a novel real-time, high quality, and training-free approach to enhance the efficiency of DiT-based video generation. PAB reduces redundancy through pyramid-style broadcasting by exploiting the U-shaped attention pattern in the diffusion process. Moreover, the broadcast sequence parallelization method significantly improves distributed inference efficiency. Overall, PAB enables real-time, high-quality video generation of up to 720p video, consistently outperforming baseline methods across various models. We believe that PAB provides a simple yet effective foundation for advancing future research and practical applications in video generation.
@@
Our approach shows promise but has some limitations. PAB’s performance may vary depending on the input data’s complexity, especially with dynamic scenes. The fixed broadcast strategy might not work best for all video types and tasks. Also, we only focused on reducing redundancy in attention mechanisms, not other parts of the model like Feed-Forward Networks. Future work could explore ways to make PAB more flexible and effective across different applications, such as developing adaptive strategies and expanding redundancy reduction to other model components.
Supplementary Material
Appendix A Experiment setup
@@ A.1 Models
As we focus on DiT-based video generation, three popular state-of-the-art open-source DiT-based video generation models are selected in the evaluation, including Open-Sora (Zheng et al., 2024), Open-Sora-Plan (Lab & etc., 2024), and Latte (Ma et al., 2024a). Open-Sora-Plan (Lab & etc., 2024) utilizes CausalVideoVAE to compress visual representations and DiT with the 3D full attention module. Open-Sora (Zheng et al., 2024) combines 2D-VAE with 3D-VAE for better video compression and uses an SD-DiT block in the diffusion process. Latte (Ma et al., 2024a) uses spatial Transformer blocks and temporal Transformer blocks to capture video information in the diffusion process.
A.2 Metrics
@@
VBench (Huang et al., 2024)
: VBench is a benchmark suite designed for evaluating video generative models, which uses a hierarchical approach to break down ’video generation quality’ into various specific, well-defined dimensions. Specifically, VBench comprises 16 dimensions in video generation, including Subject Consistency, Background Consistency, Temporal Flickering, Motion Smoothness, Dynamic Degree, Aesthetic Quality, Imaging Quality, Object Class, Multiple Objects, Human Action, Color, Spatial Relationship, Scene, Appearance Style, Temporal Style, Overall Consistency. In experiments, we adopt the VBench evaluation framework and utilize the official code to apply weighted scores to assess generation quality.
@@
: a widely used metric for measuring the quality of reconstruction in image processing. PSNR is defined as:
@@
Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018)
: a metric designed to capture perceptual similarity between images more effectively than pixel-based measures. LPIPS is based on deep learning models that learn to predict perceptual similarity by training on large datasets. It measures the distance between features extracted from pre-trained deep networks. The LPIPS score is computed as:
@@
Structural Similarity Index Measure (SSIM) (Wang & Bovik, 2002)
: the similarity between two images by considering changes in structural information, luminance, and contrast. SSIM is computed as:
@@
A.3 Baselines
We employ $\Delta$-DiT (Chen et al., 2024b) and T-GATE (Zhang et al., 2024), which are cache-based methods as baselines in the evaluation.
$\Delta$-DiT Diffusion steps b k Block range
Open-Sora-Plan 150 148 2 [0, 2]
Open-Sora 30 25 2 [0, 5]
Latte 50 48 2 [0, 5]
Table 5: Configuration of $\Delta$-DiT. $b$ represents the gate step of two stages and k is the cache interval. Block range refers to the index of the front blocks that are skipped. ’Block range’ refers to the specific indices of the blocks in the DiT-based video generation model that are skipped during the process. For example, ’Block range’ [0, 2] means that the first three blocks in the model block 0, block 1, and block 2—are skipped.
$\Delta$-DiT (Chen et al., 2024b) uses the offset of hidden states(the deviations between feature maps) rather than the feature maps themselves. $\Delta$-DiT is applied to the back blocks in the DiT during the early outline generation stage of the diffusion model and on front blocks during the detail generation stage. The stage is bounded by a hyperparameter $b$, and the cache interval is $k$. Since the source code for $\Delta$-DiT is not publicly available, we implement the baseline based on the methods in the paper. Additionally, we selected the parameters based on experimental results on video generation models. We only jump the computation of the front blocks during the Outline Generation stage. The detailed configuration is shown in Table 5.
T-GATE Diffusion steps m k
Open-Sora-Plan 150 90 6
Open-Sora 30 12 2
Latte 50 20 2

### v2

@@ 3.4 Scaling ability
To evaluate our method’s scalability, we conduct a series of experiments. In each experiment, we apply PAB246 (the best quality, but less speedup) to Open-Sora as our baseline configuration, change only the video sizes, parallel method and GPU numbers.
Table 4: Communication and latency of different sequence parallel methods on 8 NVIDIA H100 GPUs. original refers to our method on single GPU. comm. represents the total communication volume to generate a 8s 480p video.
@@
We compare the scaling efficiency with and without our method using 8 GPUs in Table 3.4 for four sequence parallelism methods including Megatron-SP (Korthikanti et al., 2023), DS-Ulysses (Jacobs et al., 2023) and DSP (Zhao et al., 2024). Our broadcast sequence parallel is implemented based on DSP, and is also adaptable to other methods. The results demonstrate that: i) PAB significantly reduces communication volume for all sequence parallelism methods. Furthermore, our method achieves the lowest communication cost compared to other techniques, and achieving near-linear scaling on 8 GPUs. With a larger temporal broadcast range, it can yield even greater performance improvements. ii) Applying sequence parallelism alone is insufficient for optimal performance because of the significant communication overhead across multiple devices.
@@
We evaluate our method’s speed in terms of FPS on 8 and 16 devices. Since in inference, the batch size of diffusion is often 2 because of CFG. Therefore, we split the batch first and apply sequence parallelism to each batch; in this way, PAB can extend to 16 devices with almost linear acceleration. As shown in Figure 9, we can achieve real-time with very high FPS video generation for 480p videos on 8 devices and even for 720p on 16 devices. Note that with acceleration techniques like Torch Compiler (Ansel et al., 2024), we are able to achieve even better speed.
@@
To further investigate how our method achieves such significant speedup, we provide a breakdown of the time consumption for various components, as shown in Figure 11. The analysis reveals that the attention calculation itself does not consume a large portion of time because the sequence length for attention will be much shorter if we do attention separately for each dimension. However, attention-related operations, such as normalization and projection, are considerably more time-consuming than the attention mechanism itself, which mainly contribute to our speedup.
@@ 3.5 Visualization
As shown in Figure 12, we visualize the video results generated by our method compared to the original model. W we employ the highest quality strategy for each model. The visualized results demonstrate that our method maintains the original quality and details.
@@ 4.1 Video Generation
Early approaches of video generation primarily leveraged GANs (Goodfellow et al., 2014), VA-VAE (Van Den Oord et al., 2017), autoregressive Transformer (Rakhimov et al., 2020) and convolution-based diffusion models (Ho et al., 2022b). Recently Video generation has seen remarkable progress driven by diffusion models, which iteratively refine noisy inputs to generate high-fidelity video frames (Ho et al., 2022b; An et al., 2023; Esser et al., 2023; Chen et al., 2024a). While many works focus on conv-based diffusions (Ho et al., 2020; Harvey et al., 2022; Singer et al., 2022; Ho et al., 2022a; Luo et al., 2023; Wang et al., 2023b; Zhang et al., 2023) and achieve good results, researchers begin to explore Transformer-based diffusion models for video generation (Zheng et al., 2024; Lab & etc., 2024; Ma et al., 2024a) because of scalability and efficiency (Peebles & Xie, 2023).
@@ 4.2 Diffusion Model Acceleration
Scheduler.
Reducing the sampling steps with schedulers has been explored through methods such as DDIM (Song et al., 2020), which enables fewer sampling steps without compromising generation quality. Other works also explore efficient solver of ODE or SDE (Song et al., 2021; Jolicoeur-Martineau et al., 2021; Lu et al., 2022; Karras et al., 2022; Lu et al., 2023).
Compression.
Researchers aimed at reducing the workload and inference time at each sampling step, including distillation (Salimans & Ho, 2022; Li et al., 2023d), quantization (Li et al., 2023c; He et al., 2023; So et al., 2023a; Shang et al., 2023) and joint optimization (Li et al., 2023a; Liu et al., 2023). However, these methods demand significant resources and data for training, making them impractical especially for large-scale pre-trained models.
Caching.
Recently, researchers revisited the concept of caching (Smith, 1982) in video generation for training-free acceleration. Some works (Ma et al., 2024c; Li et al., 2023b; Wimbauer et al., 2024; So et al., 2023b) reuse high-level features in U-Net structures while only updating low-level features based on the observation. However, these convolution-based methods can not directly apply to video DiTs. For transformer architectures, T-GATE (Zhang et al., 2024) introduce caching different attention at different stages, while $\Delta$-DiT (Chen et al., 2024c) propose to cache feature offsets of DiT blocks. Nevertheless, neither approach effectively addresses the unique attention features present in video DiTs, resulting in suboptimal performance.
Parallelism.
Sequence parallelism techniques (Korthikanti et al., 2023; Jacobs et al., 2023; Zhao et al., 2024) have been proposed to reduce generation latency through distributed inference. However, these methods introduce additional communication costs. To address this issue, some works (Li et al., 2024; Wang et al., 2024) leverage convolutional features to reduce communication overhead. Nevertheless, these approaches are still limited to convolutions.
@@ 5 Discussion and Conclusion
In this work, we introduce Pyramid Attention Broadcast (PAB), a real-time, high quality, and training-free approach to enhance the efficiency of DiT-based video generation. PAB reduces attention redundancy through pyramid-style broadcasting by exploiting the U-shaped attention pattern in the diffusion process. Moreover, our broadcast sequence parallel significantly improves distributed inference efficiency. Overall, PAB achieves up to 10.5$\times$ speedup with negligible quality loss and consistently outperforms baselines across various models. We believe that PAB provides a simple yet effective foundation for advancing future research and practical applications in video generation.
@@
PAB’s performance may vary depending on the input data’s complexity, especially with dynamic scenes. The fixed broadcast strategy might not work best for all video types and tasks. Also, we only focused on reducing redundancy in attention, not other parts of the model.
Future works.
Our work opens several promising avenues for future research. One key direction is extending to more video models with diverse architectures. Another area is the redundancy in MLPs remains under-explored and warrants further investigation. Furthermore, our findings suggest potential for developing more efficient attention for video generation.
Appendix
We organize our appendix as follows:
Experimental Settings:
  • Section A.1: Models
  • Section A.2: PAB generation settings
    • Section A.2.1: Attention
    • Section A.2.2: MLP
  • Section A.3: Baselines generation settings
  • Section A.4: Metrics
Additional Experimental Results and Findings:
  • Section B.1: Additional quantitative results
  • Section B.2: Findings for MLP broadcast
  • Section B.3: Results for long, complex and dynamic scenes
  • Section B.4: Breakdown of PAB’s contribution with multiple GPUs
  • Section B.5: Breakdown of time cost within attention module
  • Section B.6: Workflow comparison for broadcasting different objects
  • Section B.8: Extension to Text-to-Image model
  • Section B.7: Various metrics for evaluating redundancy
Appendix A Experiment Settings
@@ A.1 Models
As we focus on DiT-based video generation, three popular state-of-the-art open-source DiT-based video generation models are selected in the evaluation, including Open-Sora (Zheng et al., 2024), Open-Sora-Plan (Lab & etc., 2024), and Latte (Ma et al., 2024a). Open-Sora-Plan (Lab & etc., 2024) utilizes CausalVideoVAE to compress visual representations and DiT with the 3D full attention module. Open-Sora (Zheng et al., 2024) combines 2D-VAE with 3D-VAE for better video compression and uses an SD-DiT block in the diffusion process. Latte (Ma et al., 2024a) uses spatial Transformer blocks and temporal Transformer blocks to capture video information in the diffusion process. The inference configs of three models are shown in Table 5, which strictly follow the official settings.
Table 5: The inference config of three models.
model scheduler inference steps
Open-Sora RFLOW 30
Open-Sora-Plan PSNR 150
Latte DDIM 50
A.2 PAB Generation Settings
A.2.1 Attention
Table 6: The attention broadcast configuration of PAB. diffusion timesteps represents the start and end diffusion timestep of the broadcast, where 1000 is the beginning and 0 is the end.
model method broadcast range diffusion timesteps
spatial temporal cross
Open-Sora PAB246 2 4 6 [930-450]
PAB357 3 5 7
PAB579 5 7 9
Open-Sora-Plan PAB246 2 4 6 [850-100]
PAB357 3 5 7
PAB579 5 7 9
Latte PAB235 2 3 5 [800-100]
PAB347 3 4 7
PAB469 4 6 9
In Table 6 we demonstrate the detailed settings of attention broadcast in experiments.
A.2.2 MLP
As demonstrated in Section 2.3 and Figure 4(c), the attention outputs exhibit low difference across approximately 70% of the diffusion steps within the middle segment. Spatial attention shows the highest variance, followed by temporal attention and, finally, cross-attention. Empirically, we perform a similar analysis on the MLP module to investigate whether it also involves redundant computations during the diffusion process.
In our current evaluation experiments, we select the skippable MLP modules for each model through empirical analysis in Appendix B.2. We show our detailed configuration for MLP modules in Table 7.
Table 7: The MLP broadcast configuration of PAB. diffusion timesteps represents the starting diffusion timestep of the broadcast and the Block indicates the index of the broadcast block.
model diffusion timesteps block broadcast range
Open-Sora [864, 788, 676] [0, 1, 2, 3, 4] 2
Open-Sora-Plan [738, 714, 690, 666, 642, [0, 1, 2, 3, 4, 5, 6] 2
618, 594, 570, 546, 522,
498, 474, 450, 426]
Latte [720, 640, 560, 480, 400] [0, 1, 2, 3, 4] 2
A.3 Baselines Generation Settings
We employ $\Delta$-DiT (Chen et al., 2024c) and T-GATE (Zhang et al., 2024), which are cache-based methods as baselines in the evaluation.
Table 8: Configuration of $\Delta$-DiT. $b$ represents the gate step of two stages and k is the cache interval. Block range refers to the index of the front blocks that are skipped. Block range refers to the specific indices of the blocks in the DiT-based video generation model that are skipped during the process. For example, Block range [0, 2] means that the first three blocks in the model block 0, block 1, and block 2—are skipped.
$\Delta$-DiT diffusion steps b k block range
Open-Sora 30 25 2 [0, 5]
Open-Sora-Plan 150 148 2 [0, 1]
Latte 50 48 2 [0, 2]
$\Delta$-DiT (Chen et al., 2024c) uses the offset of hidden states (the deviations between feature maps) rather than the feature maps themselves. $\Delta$-DiT is applied to the back blocks in the DiT during the early outline generation stage of the diffusion model and on front blocks during the detail generation stage. The stage is bounded by a hyperparameter $b$, and the cache interval is $k$. Since the source code for $\Delta$-DiT is not publicly available, we implement the baseline based on the methods in the paper. Additionally, we selected the parameters based on experimental results on video generation models. We only jump the computation of the front blocks during the Outline Generation stage. The detailed configuration is shown in Table 8.
T-GATE diffusion steps m k
Open-Sora 30 12 2
Open-Sora-Plan 150 90 3
Latte 50 20 2
A.4 Metrics
@@
VBench.
VBench (Huang et al., 2024) is a benchmark suite designed for evaluating video generative models, which uses a hierarchical approach to break down ’video generation quality’ into various specific, well-defined dimensions. Specifically, VBench comprises 16 dimensions in video generation, including Subject Consistency, Background Consistency, Temporal Flickering, Motion Smoothness, Dynamic Degree, Aesthetic Quality, Imaging Quality, Object Class, Multiple Objects, Human Action, Color, Spatial Relationship, Scene, Appearance Style, Temporal Style, Overall Consistency. In experiments, we adopt the VBench evaluation framework and utilize the official code to apply weighted scores to assess generation quality.
@@
. PSNR is a widely used metric for measuring the quality of reconstruction in image processing. It is defined as:
@@
Learned Perceptual Image Patch Similarity (LPIPS).
LPIPS (Zhang et al., 2018) is a metric designed to capture perceptual similarity between images more effectively than pixel-based measures. It is based on deep learning models that learn to predict perceptual similarity by training on large datasets. It measures the distance between features extracted from pre-trained deep networks. The LPIPS score is computed as:
@@
Structural Similarity Index Measure (SSIM)
. SSIM (Wang & Bovik, 2002) evaluate the similarity between two images by considering changes in structural information, luminance, and contrast. SSIM is computed as:
@@
Appendix B Additional Experimental Results and Findings
B.1 Additional quantitative results.
model method PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$
w/ g.t. w/ org. w/ g.t. w/ org. w/ g.t. w/ org.
Open-Sora original 8.62 0.7582 0.3506
$\Delta$-DiT 9.44 12.01 0.7397 0.5263 0.3387 0.4676
T-GATE 8.38 14.22 0.7658 0.3951 0.3811 0.6286
PAB246 8.69 26.53 0.7652 0.1001 0.3606 0.8635
PAB357 8.79 24.12 0.7719 0.1597 0.3695 0.8133
PAB579 8.84 22.48 0.7821 0.2129 0.3741 0.7745
Open-Sora- Plan original 8.32 0.7701 0.2619
$\Delta$-DiT 7.88 12.26 0.7719 0.5572 0.1884 0.3865
T-GATE 8.39 13.60 0.7734 0.4750 0.2436 0.4544
PAB246 8.65 19.84 0.7653 0.2575 0.2759 0.6847
PAB357 8.87 17.39 0.7637 0.3814 0.2766 0.5767
PAB579 9.20 16.06 0.7610 0.4905 0.3025 0.4831
Latte original 8.83 0.7670 0.3008
$\Delta$-DiT 7.09 9.64 0.8071 0.7787 0.0741 0.1567
T-GATE 9.27 19.13 0.7655 0.2585 0.3202 0.6416
PAB235 9.94 19.18 0.7743 0.2667 0.3742 0.6461
PAB347 10.38 17.49 0.7775 0.3577 0.4032 0.5813
PAB469 10.60 16.76 0.7832 0.3934 0.4190 0.5619
Table 10: Quality results on webvid. Latency and speedup are calculated on one GPU. PABαβγ denotes broadcast ranges of spatial ($\alpha$), temporal ($\beta$), and cross ($\gamma$) attentions. Video generation specifications: Open-Sora (2s, 480p), Open-Sora-Plan (2.7s, 512x512), Latte (2s, 512x512). w/ g.t. indicates evaluating the metrics based on the ground-truth videos, while w/ org. means with the original methods’ outputs.
In Section 3.2, we present results only based on Vbench prompts. To further evaluate the efficacy of our method, we expand our analysis using a subset of 1000 videos from WebVid (Bain et al., 2021), a large-scale text-video dataset sourced from stock footage websites. We apply PAB to this subset, assessing its performance across three models and four metrics. The results of this additional experimentation are summarized in Table 10.
B.2 Findings for MLP broadcast.
{subcaptionblock}
0.45 Refer to caption {subcaptionblock}0.45 Refer to caption
Figure 13: Average MSE of spatial MLP
Figure 14: Average MSE of temporal MLP
{subcaptionblock}
0.45 Refer to caption {subcaptionblock}0.45 Refer to caption
Figure 15: MSE of different layers of spatial MLP
Figure 16: MSE of different layers of temporal MLP
Figure 17: Quantitative analysis of MLP module differences in Latte by mean squared error (MSE) of the MLP output across continuous time steps. In Figures (a) and (c), we present the results for the spatial MLP, while Figures (b) and (d) illustrate the outcomes for the temporal MLP. Additionally, in Figure (a) and (b), we show the average MSE across all layers. In Figures (c) and (d), we select the block 0, 8, 16, and 24 to illustrate the characteristics of MLPs across different layers.
{subcaptionblock}
0.45 Refer to caption {subcaptionblock}0.45 Refer to caption
Figure 18: Average MSE of spatial MLP
Figure 19: Average MSE of temporal MLP
{subcaptionblock}
0.45 Refer to caption {subcaptionblock}0.45 Refer to caption
Figure 20: MSE of different layers of spatial MLP
Figure 21: MSE of different layers of temporal MLP
Figure 22: Quantitative analysis of MLP module differences in Open-Sora by mean squared error (MSE) of the MLP output across continuous time steps. Figures (a) and (c) present the results for the spatial MLP, while Figures (b) and (d) show the outcomes for the temporal MLP. Figures (a) and (b) display the average MSE across all layers, and Figures (c) and (d) examine block 0, 8, 16, and 24 to showcase the MLP characteristics across different layers.
{subcaptionblock}
0.45 Refer to caption {subcaptionblock}0.45 Refer to caption
Figure 23: Average MSE of spatial MLP
Figure 24: Average MSE of temporal MLP
{subcaptionblock}
0.45 Refer to caption {subcaptionblock}0.45 Refer to caption
Figure 25: MSE of different layers of spatial MLP
Figure 26: MSE of different layers of temporal MLP
Figure 27: Quantitative analysis of MLP module differences in Open-Sora-Plan by mean squared error (MSE) of the MLP output across continuous time steps. In Figures (a) and (c), the results for the spatial MLP are shown, while Figures (b) and (d) show the results for the temporal MLP. Figures (a) and (b) display the average MSE across all layers, and Figures (c) and (d) highlight block 0, 8, 16, and 24 to illustrate the MLP behavior across different layers.
We present a quantitative analysis of the FFN output differences in Latte, Open-Sora, and Open-Sora-Plan, using mean squared error (MSE) as the evaluation metric in Figure 17,  22 and  27.
We observe that during the intermediate stages of diffusion, the MSE exhibits a periodic spiking pattern, where local maxima occurs at specific timesteps, followed by consistently low values in subsequent timesteps. Therefore, we can retain the MLP output at the peak and reuse it during the following low-value timesteps. Additionally, by analyzing the FFN modules across different blocks in Figure, we found that the output differences in the lower layers’ MLPs are relatively small, while those in the upper layers’ MLPs are significantly larger. Based on these findings, we empirically selected MLP modules to broadcast and corresponding broadcast ranges for each model, including Latte, Open-Sora, and Open-Sora-Plan.
B.3 Results for long, complex and dynamic scenes.
In this section, we evaluate the quantitative and qualitative results for PAB when dealing with long, complex and dynamic scenes.
Table 11: Quantitative results of Open-Sora (16s 480p) on subset dimensions of Vbench datasets for long, complex and dynamic scenes.
method human overall imaging aesthetic dynamic motion subject total
action consistency quality quality degree smoothness consistency
original 92.67 73.65 61.40 56.59 21.07 96.43 90.26 74.54
PAB246 91.33 73.43 60.18 56.24 19.91 97.05 90.08 73.98
PAB357 89.33 72.53 58.17 54.86 19.45 96.40 88.35 72.58
PAB579 88.33 72.36 57.92 54.63 18.05 96.50 88.32 72.15
Quantitative results.
For model settings, we specifically use Open-Sora to generate videos of 16 seconds duration. This longer duration purposefully challenges our method with more complex and dynamic scenes. Open-Sora is the only model used as other models are restricted to fixed short lengths.
For dataset, from VBench’s comprehensive 16-dimensional evaluation metrics, we strategically select 7 categories that best assess complex and dynamic scenes. Total performance scores are calculated based exclusively on these 7 categories to provide focused evaluation of complex and dynamic capabilities.
As shown in Table 11, we specifically test our method under the most challenging conditions by using the longest videos and selecting the more complex tasks in the dataset. The results show that PAB performs consistently well, with PAB246 showing comparable performance.
What’s particularly encouraging is that PAB maintains good scores even in the most difficult dimensions we tested, like human action and dynamic degree. This shows that our model stays reliable even under demanding conditions.
Qualitative results.
As shown in Figure 28, our method demonstrates robust performance in processing dynamic, complex scenes while maintaining high-quality output.
Figure 28: Qualitative results of Open-Sora (16s 480p) on subset dimensions of Vbench datasets for long, complex and dynamic scenes.
B.4 Breakdown of PAB’s contribution with multiple GPUs
As shown in Table B.4, we evaluate the independent contribution of PAB from computation and communication and come with the following conclusions:
  • With PAB’s computation speedup (save computation by attention broadcast), the latency is further reduced by 24% compared with DSP only.
  • With PAB’s communication speedup (can save all communication cost when temporal attention is skipped), the latency can be further reduced by 5.0% compared with computation speedup only.
  • Since we only evaluate based on PAB246 (better quality but less speedup), PAB is able achieve more speedup if using more aggressive strategies.
Table 12: Breakdown of PAB’s contribution with multiple GPUs using Open-Sora (8s 480p).
method latency (s)
original (1 gpu) 96.90
DSP (8 gpus) 11.53
DSP + PAB (with computation speedup) (8 gpus) 9.29
DSP + PAB (with computation and communication speedup) (8 gpus) 8.85
B.5 Breakdown of time cost within attention module
As shown in Figure 11, the attention operation takes only a small proportion of time in attention module for 2s 480p Open-Sora. In this section, we further investigate what the main cost in attention module.
As shown in Table 13;14;15, our findings show that attention operation isn’t actually the main thing slowing down the model. The real bottleneck comes from other parts - specifically layernorm and positional embedding. Even though these operations have fewer calculations and FLOPs, they run much slower in practice. Because modern GPUs are built to handle big matrix calculations super efficiently, but they struggle with operations that work on one element at a time, which is exactly what LayerNorm and positional embedding do.
Table 13: Breakdown of time cost in spatial attention.
time layernorm1 mask modulate qkv proj layernorm2 o proj attn reshape
absolute (ms) 1.132 0.149 0.616 0.473 2.160 0.176 1.595 0.312
normalized 17.1% 2.2% 9.3% 7.2% 32.7% 2.7% 24.1% 4.7%
Table 14: Breakdown of time cost in temporal attention.
time layernorm1 mask modulate qkv proj layernorm2 pos emb o proj attn reshape
absolute (ms) 1.126 0.150 0.616 0.477 2.154 2.610 0.176 0.896 0.314
normalized 13.1% 1.8% 7.2% 5.6% 25.3% 30.7% 2.1% 10.6% 3.6%
Table 15: Breakdown of time cost in cross attention.
time qkv proj attn o proj reshape
absolute (ms) 0.771 0.362 0.176 0.912
normalized 34.8% 16.2% 7.9% 41.1%
B.6 Workflow comparison for broadcasting different objects
In Table 3.3, we demonstrate the efficiency of broadcasting different objects. In this section, we further demonstrate why there will be such difference:
  • Broadcasting attention outputs enables us to bypass all intermediate computations within the attention module (including layer normalization, positional embedding, and qkvo projections) while maintaining compatibility with efficient attention kernels such as FlashAttention (we enable FlashAttention in all experiments by default to be closer to real-world usage).
  • But broadcasting attention scores still requires partial computation in the attention module (e.g., attention calculation and linear projection). Its performance may even degrade below baseline due to incompatibility with FlashAttention.
To be more clear, here are the workflows in attention module for different broadcast strategies:
  • original:
    $x\to q,k,v=proj(x)\to q,k=pos\_emb(layer\_norm(q,k))\to o=attn(q,k,v)\to o=proj(o)$
  • attention score (cannot use FlashAttention because we need attention score explicitly):
    $x\to v=proj(x)\to o=attn(broadcast\_score,v)\to o=proj(o)$
  • attention outputs:
    $o=broadcast\_outputs$
B.7 Various metrics for evaluating redundancy
We evaluate different metrics for measuring redundancy as shown in Figure 29.
Figure 29: Various metrics for evaluating redundancy.
B.8 Extension to Text-to-Image model
PAB also has the potential to extend to Text-to-Image model like FLUX. In this section, we demonstrate our speedup, qualitative and quantitative results on FLUX.
Figure 30: Qualitative results of FLUX with PAB55.
Table 16: Speedup of FLUX. PABαγ denotes broadcast ranges of spatial ($\alpha$) and cross ($\gamma$) attentions.
method latency (s)
original 13.8
PAB55 7.8
As shown in Table 16, we can achieve 1.77$\times$ speedup compared with original method. We choose PAB55 because it offers best balance between speedup and image quality. Note that this is only ran on single GPU.
As shown in Figure 30, we visualize the quantitative results on FLUX. Our method can achieve comparable results compared with baseline.